OpenStudy (loser66):

A series circuit has a capacitor of 0.25*10^-6 farad, a resistor of 5* 10^3 ohms, and an inductor of 1 henry. The initial charge on the capacitor is 0. If a 12volt battery is connected to the circuit and the circuit is closed at t=0, determine the charge on the capacitor at t = 0.001sec, at t= 0.01 sec and at any time t. Also, determine the limiting charge as t --> infinitive Please, help

OpenStudy (loser66):

@druminjosh

OpenStudy (anonymous):

word problem wonderful

OpenStudy (loser66):

No, it killed me, it was my test. I knew nothing about it,

OpenStudy (anonymous):

you know its a second order ODE yes?

OpenStudy (loser66):

yes, I had equation of this

OpenStudy (anonymous):

and you can solve it?

OpenStudy (loser66):

I just copied from the note and solved it. I never took physics before, know nothing.

OpenStudy (anonymous):

\[L \frac{ d^2q }{ dt^2 }+R \frac{ dq }{ dt }+\frac{ 1 }{ C }\frac{ dq }{ dt }=E(t)\]

OpenStudy (loser66):

yes,

OpenStudy (anonymous):

ok so we fill in the constants from what's given, yes?

OpenStudy (loser66):

yes,

OpenStudy (anonymous):

ok let me do that see what i get... i'm not sure if that 12 volt battery makes E(t) = 12.

OpenStudy (anonymous):

hang on hang on

OpenStudy (loser66):

\[Q" + 5*10^3 Q' +\frac{1}{0.25*10^{-6}}Q=12\]

OpenStudy (anonymous):

ok 10^-6 in denominator is 10^6 in numerator so i wrote like this: \[Q''+5000Q'+25000Q=12\]

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

wait no hang on grr

OpenStudy (anonymous):

should be 4000000Q yes?

OpenStudy (anonymous):

\[Q''+5000Q'+4000000Q=12; Q(0)=0\]

OpenStudy (loser66):

yes, but I let the whole thing there to solve

OpenStudy (anonymous):

but there's nothing about what happens at Q'(0) hrmm

OpenStudy (anonymous):

So \[y_c = C_1e^{-1000t}+C_2e^{-4000t}\] yes?

OpenStudy (loser66):

:(./......... :( hu huhu... I got it wrong,!!! My bad, I was too nervous to solve that second order differential equation.

OpenStudy (anonymous):

but not sure about the constants because i'm not certain about what happens to the first derivative at 0

OpenStudy (anonymous):

oh they're both at 0 ... the circuit is closed at 0 and the initial charge on the capacitor is 0.

OpenStudy (anonymous):

you can use variation of parameters to solve it LOL

OpenStudy (loser66):

Ok, you help me solve all of them . I need the whole thing to study. I still have a chance to make up the test.

OpenStudy (anonymous):

i was kidding about variation of parameters :)

OpenStudy (anonymous):

can use coefficients, \[y_p=At^2+Bt+C\] yes?

OpenStudy (loser66):

It took me 3 days to pass the depress. Now it is back to me.

OpenStudy (loser66):

nope, the yp is A only, it's a constant

OpenStudy (loser66):

Am I right?

OpenStudy (anonymous):

well the second derivative of a quadratic function can be 12, i thought you need the bt and c terms too

OpenStudy (loser66):

there is something wrong at my computer, the letter jumps up and down.

OpenStudy (loser66):

why? My prof taught me that, when the RHS is a constant, the "guessing" of partial solution is an A only

OpenStudy (anonymous):

most likely b and c will be 0 ok assume its just At^2

OpenStudy (loser66):

Josh, I don't get!!! whyyyyy??? if it is a polynomial, I can understand yours. This case is not that,

OpenStudy (anonymous):

i think you have to use at^2+bt+c but hang on i'm working on it ok bro?

OpenStudy (anonymous):

its possible that the derivative of at^2+bt+c is 12

OpenStudy (anonymous):

2nd derivative

OpenStudy (loser66):

yes,

OpenStudy (anonymous):

actually i think its just A but hang on, i don't do these every day lol

OpenStudy (loser66):

hehehe,, take time, bro

OpenStudy (anonymous):

yes the particular solution is of the form y_p = A i am pretty sure

OpenStudy (loser66):

hey, you confused me. Actually, what is the form?

OpenStudy (anonymous):

i have a different problem in a book, y"+3y'+2y= 6 and the particular solution is 3 so... i think its just A... the right hand side is a polynomial of degree 0 so the particular solution is a polynomial with degree 0

OpenStudy (loser66):

hahahaa... we meet there.

OpenStudy (anonymous):

we could do variation of parameters if we're unsure about the method of coefficients

OpenStudy (anonymous):

lets try A, if y_p = A, then both derivatives are 0 ... so we get that 4000000A = 12, so A = 3e-6

OpenStudy (loser66):

yes,

OpenStudy (anonymous):

so the particular solutio is just 3e-6 yes?

OpenStudy (anonymous):

3E-6 = 3*10^-6

OpenStudy (anonymous):

3/1000000

OpenStudy (loser66):

I have question there. In non homogeneous ODE, we have to combine y homo and y partial, then put the initial condition to get the constant, right?

OpenStudy (loser66):

or separate them , solve part by part?

OpenStudy (anonymous):

you have to find c1 and c2 with the particular solution

OpenStudy (loser66):

got you

OpenStudy (loser66):

ok, thank you very much Josh

OpenStudy (anonymous):

so what you get for solution?

OpenStudy (loser66):

I forgot. I was too sad after test, didn't want to think about it. Tomorrow I have class, have to face it on. but I forgot what I was doing.

OpenStudy (anonymous):

i meant now... i gave you general solution, you just gotta find c1 and c2

OpenStudy (loser66):

hey, Josh, I need you guide me more about Fourier and Laplace when I study them. Visit this site frequently ,please

OpenStudy (anonymous):

send me an e-mail, sometimes i get pretty busy... fourier and laplace transforms are my favorite

OpenStudy (loser66):

you give me the initial condition Q(0) = 0 and Q'(0) =0 I can solve for C1 and C2. don't worry

OpenStudy (anonymous):

ok its ugly linear algebra

OpenStudy (loser66):

email? what do you mean?

OpenStudy (anonymous):

oh nm when you post my name in here, it e-mails me and it comes up on my phone. if i'm not busy i'll help out

OpenStudy (loser66):

ok, that email not this site pm, right?

OpenStudy (anonymous):

pm yes

OpenStudy (loser66):

how can you link them together? this site and your email?

OpenStudy (anonymous):

it automatically e-mails me when someone messages me or mentions me in a post i don't know how to be honest

OpenStudy (anonymous):

I think its under settings

OpenStudy (loser66):

ah, you set it in your private profile. right?

OpenStudy (anonymous):

Under Settings --> Notification Settings

OpenStudy (loser66):

I used to set it until it turned annoying. I reset. so I forgot that part. OK. I will send you if I got stuck Thanks a lot